### v1

@@ 3.2 The implementation
The challenge that we need to face is now the computation of $\tilde{T}^{\mathrm{AB}}_{n;\,\ell}(k)$, the integrals of the two Bessel functions and the Chebyshev polynomials, with the appropriate $k$-factor for each case. Those integrals need to be computed with high accuracy as they constitute the central step of the algorithm. We performed this integration using the Clenshaw-Curtis quadrature rule (Clenshaw & Curtis, 1960) and $N=2^{15}+1$ points in the interval $k_{\text{min}}-k_{\text{max}}$. The value for $k_{\text{min}}$ is $3.571\times 10^{-4}$, and $k_{\text{max}}$ is $15.385$; the choice of those values will be motivated in Section 4. $N=2^{15}+1$ is the value that provides best balance between accuracy and performance for our purposes. In this setting, all the $\tilde{T}^{\mathrm{AB}}_{n;\,\ell}(k)$ can be computed on a laptop in $\approx 1$-$2$ hours with $n_{\text{max}}$, the number of Chebyshev polynomials in the power spectrum approximation, fixed to $120$. How the accuracy scales with respect to the parameter $N$ is discussed in Appendix C.
As it is the central part of the algorithm, we highlight that, whichever the choice of $N$ and $n_{\text{max}}$, this computation is only to be performed once. It is also worth mentioning that the projected matter densities $w_{\ell}^{\mathrm{AB}}(\chi,R\chi)$ that we computed were validated using two independent and well-established codes: TWOFAST (Grasshorn Gebhardt & Jeong, 2018) and QuadOsc (Press, 2007). TWOFAST is a Julia implementation of the FFTLog algorithm (Talman, 1978; Hamilton, 2000), that will be discussed below. QuadOsc solves the problem of the highly oscillatory integrand by integrating between its zero-crossings and successively adding up the contributions. Our evaluation of the integral was found to be in excellent agreement with both codes, confirming the validity of our new approach.
@@
The last thing to mention concerns the matter power spectrum. The method we developed treats the linear, $P_{\text{lin}}(k)$, and non-linear, $P_{\delta}(k)$, matter power spectrum as two separate components. In particular, the splitting we do is the following:
$P_{\delta}(k,\chi_{1},\chi_{2})=P_{\text{lin}}(k,\chi_{1},\chi_{2})+[P_{\delta}-P_{\text{lin}}](k,\chi_{1},\chi_{2}).$ (29)
@@
$C^{\text{tot}}(\ell)=C_{\text{lin}}(\ell)+C_{\delta}^{\text{limb}}(\ell)-C_{\text{lin}}^{\text{limb}}(\ell).$ (30)
@@ 4.1 Challenge set-up
$\Delta\chi^{2}\equiv\sum_{b}N_{b}\text{Tr}(\Sigma_{b}^{-1}\Delta C_{b})^{2}.$ (31)
$\Delta C_{\ell}$ is the difference between the benchmark prediction and the prediction from a given non-Limber method. $N_{b}$ is the effective number of modes in that specific bandpower, defined as:
$N_{b}=\frac{f_{\text{sky}}}{2}\sum_{\ell\in b}(2\ell+1).$ (32)
For $f_{\text{sky}}$ we are using the sky fraction of LSST, i.e. $0.4$. Finally, $\Sigma$ is defined as:
$\Sigma^{\mathrm{AB}}_{ij}(\ell)=\sqrt{\frac{2}{(2\ell+1)\Delta\ell f_{sky}}}\left[C_{ij}^{\mathrm{AB}}(\ell)+N^{\mathrm{AB}}_{ij}(\ell)\right].$ (33)
$\Sigma$ is a $15\times 15$ matrix containing the benchmark $C_{\ell}$’s in bandpower $b$ and the LSST shape-noise and shot-noise terms ($\epsilon_{s}^{2}/n_{s}$, and $1/n_{l}$ with $\epsilon_{s}=0.28$, $n_{s}=27$ source galaxies per arcmin2, and $n_{l}=40$ lens galaxies per arcmin2). It can be shown (Hamimeche & Lewis, 2008; Carron, 2013) that the definition of $\Delta\chi^{2}$ in Eq. (31) is equivalent to the standard $\Delta\chi^{2}=\Delta\mathbf{x}^{T}\text{Cov}^{-1}\Delta\mathbf{x}$, Cov being the Gaussian covariance of the angular power spectra. The required accuracy is $\Delta\chi^{2}<0.2$ for $\ell\in(2,200)$.
@@ 4.2.1 FKEM (CosmoLike)
$\begin{split}P_{\delta}(k,z(\chi_{1}),z(\chi_{2}))&=D(\chi_{1})D(\chi_{2})P_{\text{lin}}(k,z=0)\\ &\quad+[P_{\delta}-P_{\text{lin}}](k,\chi_{1},\chi_{2})\end{split}$ (34)
@@
When choosing $\nu_{n}=2\pi i\cdot n/N+\eta$, the power law becomes $x^{\nu}=\exp{(2\pi i\cdot n/N\cdot\log{x})}\cdot x^{\eta}$, which is a Fourier Transform in the variable $\log{x}$. The integrals of the product of Bessel functions and power laws have analytical solutions, which can be efficiently computed as it is possible to use the FFT algorithm thanks to the choice of $\nu_{n}$. For a more detailed description of FKEM and the FFTLog algorithm, see (Leonard et al., 2023; Fang et al., 2020).
@@ 7 Conclusions
The development of accurate and fast algorithms for computing cosmological quantities is of crucial importance considering the unprecedented amount and quality of data that next-generation cosmological surveys are about to bring into the field. In this paper we focused on improving the evaluation of the model for $3\times 2$ pt statistics, without the employment of the Limber approximation. We presented Blast.jl, a new algorithm to compute non-Limber angular power spectra for galaxy clustering and cosmic shear samples, whose core idea is a decomposition of the power spectrum into the complete and orthonormal basis of the Chebyshev polynomials. By expressing the 3D matter power spectrum in terms of this basis, the most computationally challenging part of the integral can be effectively isolated, allowing for it to be pre-computed and stored. This significantly reduces the complexity of the overall computation, making it more efficient. The performance of our new algorithms was assessed in a fair and robust way, according to the standards of the N5K challenge, a non-Limber integration challenge launched in 2020 by LSST DESC (Ivezic et al., 2019) that allowed us to test Blast.jl against the state-of-the-art non-Limber algorithms. In terms of speed of computation at the required accuracy, Blast.jl outperformed FKEM, the winner of the challenge. In particular, our algorithm is $10$-15×15$\times$15 × faster in the fiducial scenario. Blast.jl runtime also proved to be basically independent of the number of tomographic redshift bins in the analysis and is also insensible to the accuracy requirements, making Blast.jl the preferred method in those scenarios. Another scaling we explored is the accuracy as a function of the width of the redshift bins: while the performance is still comparable to that of the other challenge entries, matter is still the least sensitive algorithm with respect to this specific hyper-parameter, as it keeps essentially the same accuracy when the bin width is halved. Working on improving the performance of our algorithm in this scenario is reserved for future work. Nonetheless, the tests performed in this work show that increasing the number of integration points strongly impacts Blast.jl’s accuracy, giving a possible solution to this behaviour. We also highlight that, unlike FKEM, Blast.jl does not assume a scale-independent growth factor, so it is possible to employ it in cosmologies with massive neutrinos or modified gravity scenarios. In that sense, the algorithm we really want to compare to is matter. With respect to that entry, the speed in the fiducial scenario improves by over 2 orders of magnitude. To the best of our knowledge, the fastest non-Limber integration algorithm is the one presented in this paper; actually, only the emulation of the whole $3\times 2$ pt statistic would result in a shorter runtime (Zhong et al., 2024; Saraivanov et al., 2024).
@@
Future work will focus on adding other effects to the $C_{\ell}$’s, in particular redshift space distortions, magnification bias and relativistic corrections like the integrated Sachs-Wolfe effect. Those will enable the use of Blast.jl for a real cosmological analysis. We will also work on including the CMB as an alternative observable, allowing the user to perform galaxy-CMB lensing cross-correlation analyses. Another future development is related to the computation of the matter power spectrum; given the fast runtime of Blast.jl, it is important that this step does not become the new bottleneck of the $3\times 2$ pt computation. Currently, we are working on a matter power spectrum emulator, tailored to work in synergy with Blast.jl. Finally, Blast.jl will also include the support for automatic differentiation. Having a $3\times 2$ pt statistics code that is differentiable is crucial because it enables the use of gradient-based methods (Campagne et al., 2023; Nygaard et al., 2023; Piras et al., 2024; Bonici et al., 2024; Ruiz-Zapatero et al., 2024; Balkenhol et al., 2024; Bonici et al., 2022; Giovanetti et al., 2024). The ultimate goal is to perform cosmological parameter inference using $3\times 2$ pt statistic and a fully-differentiable likelihood that can be sampled very efficiently thanks to Blast.jl. The code will be public upon acceptance of this paper.
@@ Appendix A Choosing an effective basis for the integration
A crucial step for the successful calculation of the angular power spectra with this method is the re-parametrization of the integral from the $\chi_{1}$-$\chi_{2}$ variables to $\chi$-$R$, with $R\equiv\chi_{2}/\chi_{1}$. The projected matter densities (Eq. (24)) present a sharp feature along their diagonal ($\chi_{1}=\chi_{2}$). Capturing this feature in the $\chi_{1}$-$\chi_{2}$ reference frame would require a high number of sampling points ($>>500$), which is computationally not convenient. If we define $R$ as the ratio of the coordinates, we know that $R\in(0,1]$, and $R=1$ represents the diagonal. It is therefore easier and more intuitive to perform a finer sampling of the diagonal feature, and a coarser one outside the diagonal, where the contribution to the integral are close to zero. Fig. 6 is a visual representation of this change of basis, showing how most of the contributions to the integral come from $R>0.8$. Note also that the figure shows a low multipole $\ell=2$, for which the off-diagonal contributions are still relevant. For higher multipole orders, the feature gets sharper and sharper. This effectively shows why the Limber approximation, which accounts only for the contribution coming from the diagonal $\chi_{1}=\chi_{2}$, is more accurate at higher multipoles $\ell$.
@@
[…]
@@
[…]
@@
$\mathcal{K}_{i}^{\mathrm{A}}(\chi)=\begin{cases}K_{i}^{\mathrm{A}}(\chi)&\text{for clustering}\\ K_{i}^{\mathrm{A}}(\chi)/\chi^{2}&\text{for lensing}\end{cases}$ (38)
@@ Appendix B Assessing model errors
$\sigma_{ij}^{\mathrm{AB}}(\ell)=\sqrt{\frac{C_{ii}^{\mathrm{AA}}(\ell)C_{jj}^{\mathrm{BB}}(\ell)+(C_{ij}^{\mathrm{AB}}(\ell))^{2}}{f_{\text{sky}}(2\ell+1)}}$ (39)
where we are using $f_{\text{sky}}=0.4$, the sky fraction for LSST. In particular, we can define a metric that is often referred to as Mean Absolute Error Ratio (MAER):
$\text{MAER}^{\mathrm{AB}}_{ij}(\ell)=\frac{|C_{ij}^{\mathrm{AB}}(\ell)-C_{ij,\text{bench}}^{\mathrm{AB}}(\ell)|}{\sigma_{ij}^{\mathrm{AB}}(\ell)}$ (40)

### v2

@@ 3.2 The implementation
The challenge that we need to face is now the computation of $\tilde{T}^{\mathrm{AB}}_{n;\,\ell}(k)$, the integrals of the two Bessel functions and the Chebyshev polynomials, with the appropriate $k$-factor for each case. Those integrals need to be computed with high accuracy as they constitute the central step of the algorithm. We performed this integration using the Clenshaw-Curtis quadrature rule (Clenshaw & Curtis, 1960) and $N=2^{15}+1$ points in the interval $[k_{\mathrm{min}},k_{\mathrm{max}}]$. The value for $k_{\mathrm{min}}$ is $3.571\times 10^{-4}\,\,\mathrm{Mpc}^{-1}$, and $k_{\mathrm{max}}$ is $15.385\,\,\mathrm{Mpc}^{-1}$; the choice of those values will be motivated in Section 4. $N=2^{15}+1$ is the value that provides best balance between accuracy and performance for our purposes. In this setting, all the $\tilde{T}^{\mathrm{AB}}_{n;\,\ell}(k)$ can be computed on a laptop in $\approx 1$-$2$ hours with $n_{\mathrm{max}}$, the number of Chebyshev polynomials in the power spectrum approximation, fixed to $120$. How the accuracy scales with respect to the parameter $N$ is discussed in Appendix C.
As it is the central part of the algorithm, we highlight that, whichever the choice of $N$ and $n_{\mathrm{max}}$, this computation is only to be performed once. It is also worth mentioning that the projected matter densities $w_{\ell}^{\mathrm{AB}}(\chi,R\chi)$ that we computed were validated using two independent and well-established codes: TWOFAST (Grasshorn Gebhardt & Jeong, 2018) and QuadOsc (Press, 2007). TWOFAST is a Julia implementation of the FFTLog algorithm (Talman, 1978; Hamilton, 2000), that will be discussed below. QuadOsc solves the problem of the highly oscillatory integrand by integrating between its zero-crossings and successively adding up the contributions. Our evaluation of the integral was found to be in excellent agreement with both codes, confirming the validity of our new approach.
@@
The last thing to mention concerns the matter power spectrum. The method we developed treats the linear, $P_{\mathrm{lin}}(k)$, and non-linear, $P_{\delta}(k)$, matter power spectrum as two separate components. In particular, the splitting we do is the following:
$P_{\delta}(k,\chi_{1},\chi_{2})=P_{\mathrm{lin}}(k,\chi_{1},\chi_{2})+[P_{\delta}-P_{\mathrm{lin}}](k,\chi_{1},\chi_{2}).$ (36)
@@
$C^{\mathrm{tot}}(\ell)=C_{\mathrm{lin}}(\ell)+C_{\delta}^{\mathrm{limb}}(\ell)-C_{\mathrm{lin}}^{\mathrm{limb}}(\ell).$ (37)
@@ 4.1 Challenge set-up
$\Delta\chi^{2}\equiv\sum_{b}N_{b}\mathrm{Tr}(\Sigma_{ij}^{\mathrm{AB}}(\ell_{b})^{-1}\Delta C_{ij}^{AB}(\ell_{b}))^{2},$ (38)
where $b$ is the index identifying the bandpower considered, $\Delta C_{ij}^{AB}(\ell_{b})$ is the difference between the benchmark prediction and the prediction from a given non-Limber method. $N_{b}$ is the effective number of modes in that specific bandpower, defined as:
$N_{b}=\frac{f_{\mathrm{sky}}}{2}\sum_{\ell\in b}(2\ell+1).$ (39)
For $f_{\mathrm{sky}}$ we are using the sky fraction of LSST, i.e. $0.4$. Finally, $\Sigma^{\mathrm{AB}}_{ij}(\ell_{b})$ is defined as:
$\Sigma^{\mathrm{AB}}_{ij}(\ell_{b})=\sqrt{\frac{2}{(2\ell_{b}+1)\Delta\ell_{b}f_{sky}}}\left[C_{ij}^{\mathrm{AB}}(\ell_{b})+N^{\mathrm{AB}}_{ij}(\ell_{b})\right].$ (40)
$\Sigma^{\mathrm{AB}}_{ij}(\ell_{b})$ is a $15\times 15$ matrix containing the benchmark $C_{\ell}$’s in bandpower $b$ and the LSST shape-noise and shot-noise terms ($\epsilon_{s}^{2}/n_{s}$, and $1/n_{l}$ with $\epsilon_{s}=0.28$, $n_{s}=27$ source galaxies per arcmin2, and $n_{l}=40$ lens galaxies per arcmin2). It can be shown (Hamimeche & Lewis, 2008; Carron, 2013) that the definition of $\Delta\chi^{2}$ in Eq. (38) is equivalent to the standard $\Delta\chi^{2}=\Delta\mathbf{x}^{T}\mathrm{Cov}^{-1}\Delta\mathbf{x}$, $\mathrm{Cov}$ being the Gaussian covariance of the angular power spectra. The required accuracy is $\Delta\chi^{2}<0.2$ for $\ell\in(2,200)$.
@@ 4.2.1 FKEM (CosmoLike)
$\begin{split}P_{\delta}(k,z(\chi_{1}),z(\chi_{2}))&=D(\chi_{1})D(\chi_{2})P_{\text{lin}}(k,z=0)\\ &\quad+[P_{\delta}-P_{\text{lin}}](k,\chi_{1},\chi_{2}),\end{split}$ (41)
@@
When choosing $\nu_{n}=2\pi\mathrm{i}\cdot n/N+\eta$, the power law becomes $x^{\nu}=\exp{(2\pi\mathrm{i}\cdot n/N\cdot\log{x})}\cdot x^{\eta}$, which is a Fourier Transform in the variable $\log{x}$. The integrals of the product of Bessel functions and power laws have analytical solutions, which can be efficiently computed as it is possible to use the FFT algorithm thanks to the choice of $\nu_{n}$. For a more detailed description of FKEM and the FFTLog algorithm, see (Leonard et al., 2023; Fang et al., 2020).
@@ 7 Conclusions
The development of accurate and fast algorithms for computing cosmological quantities is of crucial importance considering the unprecedented amount and quality of data that next-generation cosmological surveys are about to bring into the field. In this paper we focused on improving the evaluation of the model for $3\times 2$ pt statistics, without the employment of the Limber approximation. We presented Blast.jl, a new algorithm to compute non-Limber angular power spectra for galaxy clustering and cosmic shear samples, whose core idea is a decomposition of the power spectrum into the complete and orthonormal basis of the Chebyshev polynomials. By expressing the $3D$ matter power spectrum in terms of this basis, the most computationally challenging part of the integral can be effectively isolated, allowing for it to be pre-computed and stored. This significantly reduces the complexity of the overall computation, making it more efficient. The performance of our new algorithms was assessed in a fair and robust way, according to the standards of the N5K challenge, a non-Limber integration challenge launched in 2020 by LSST DESC (Ivezic et al., 2019) that allowed us to test Blast.jl against the state-of-the-art non-Limber algorithms. In terms of speed of computation at the required accuracy, Blast.jl outperformed FKEM, the winner of the challenge. In particular, our algorithm is $10$-15×15$\times$15 × faster in the fiducial scenario. Blast.jl runtime also proved to be basically independent of the number of tomographic redshift bins in the analysis and is also insensitive to the accuracy requirements, making Blast.jl the preferred method in those scenarios. Another scaling we explored is the accuracy as a function of the width of the redshift bins: while the performance is still comparable to that of the other challenge entries, matter is still the least sensitive algorithm with respect to this specific hyper-parameter, as it keeps essentially the same accuracy when the bin width is halved. Working on improving the performance of our algorithm in this scenario is reserved for future work. Nonetheless, the tests performed in this work show that increasing the number of integration points strongly impacts Blast.jl’s accuracy, giving a possible solution to this behaviour. We also highlight that, unlike FKEM, Blast.jl does not assume a scale-independent growth factor, so it is possible to employ it in cosmologies with massive neutrinos or modified gravity scenarios. In that sense, the algorithm we really want to compare to is matter. With respect to that entry, the speed in the fiducial scenario improves by over 2 orders of magnitude. To the best of our knowledge, the fastest non-Limber integration algorithm is the one presented in this paper; actually, only the emulation of the whole $3\times 2$ pt statistic would result in a shorter runtime (Zhong et al., 2024; Saraivanov et al., 2024).
@@
Future work will focus on adding other effects to the $C_{\ell}$’s, in particular redshift space distortions, magnification bias and relativistic corrections like the integrated Sachs-Wolfe effect. Those will enable the use of Blast.jl for a real cosmological analysis. We will also work on including the CMB as an alternative observable, allowing the user to perform galaxy-CMB lensing cross-correlation analyses. Another future development is related to the computation of the matter power spectrum; given the fast runtime of Blast.jl, it is important that this step does not become the new bottleneck of the $3\times 2$ pt computation. Currently, we are working on a matter power spectrum emulator, tailored to work in synergy with Blast.jl. Finally, Blast.jl will also include the support for automatic differentiation. Having a $3\times 2$ pt statistics code that is differentiable is crucial because it enables the use of gradient-based methods (Campagne et al., 2023; Nygaard et al., 2023; Piras et al., 2024; Bonici et al., 2024; Ruiz-Zapatero et al., 2024; Balkenhol et al., 2024; Bonici et al., 2022; Giovanetti et al., 2024). The ultimate goal is to perform cosmological parameter inference using $3\times 2$ pt statistic and a fully-differentiable likelihood that can be sampled very efficiently thanks to Blast.jl.
@@ Appendix A Choosing an effective basis for the integration
A crucial step for the successful calculation of the angular power spectra with this method is the re-parametrization of the integral from the $\chi_{1}$-$\chi_{2}$ variables to $\chi$-$R$, with $R\equiv\chi_{2}/\chi_{1}$. The projected matter densities (Eq. (31)) present a sharp feature along their diagonal ($\chi_{1}=\chi_{2}$). Capturing this feature in the $\chi_{1}$-$\chi_{2}$ reference frame would require a high number of sampling points ($\gg 500$), which is computationally not convenient. If we define $R$ as the ratio of the coordinates, we know that $R\in(0,1]$, and $R=1$ represents the diagonal. It is therefore easier and more intuitive to perform a finer sampling of the diagonal feature, and a coarser one outside the diagonal, where the contribution to the integral are close to zero. Fig. 6 is a visual representation of this change of basis, showing how most of the contributions to the integral come from $R>0.8$. Note also that the figure shows a low multipole $\ell=2$, for which the off-diagonal contributions are still relevant. For higher multipole orders, the feature gets sharper and sharper. This effectively shows why the Limber approximation, which accounts only for the contribution coming from the diagonal $\chi_{1}=\chi_{2}$, is more accurate at higher multipoles $\ell$.
@@
$C_{ij}^{\mathrm{AB}}(\ell)$ $=\int_{0}^{\infty}\mathrm{d}\chi_{1}K_{i}^{\mathrm{A}}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}K_{j}^{\mathrm{B}}(\chi_{2})w_{\mathrm{AB}}(\chi_{1},\chi_{2})$ (43)
$=\int_{0}^{\infty}\mathrm{d}\chi_{1}K_{i}^{\mathrm{A}}(\chi_{1})\int_{0}^{\chi_{1}}\mathrm{d}\chi_{2}K_{j}^{\mathrm{B}}(\chi_{2})w_{\mathrm{AB}}(\chi_{1},\chi_{2})+\int_{0}^{\infty}\mathrm{d}\chi_{2}K_{j}^{\mathrm{B}}(\chi_{2})\int_{0}^{\chi_{2}}\mathrm{d}\chi_{1}K_{i}^{\mathrm{A}}(\chi_{1})w_{\mathrm{AB}}(\chi_{2},\chi_{1}).$
@@
$C_{ij}^{\mathrm{AB}}(\ell)=\int_{0}^{\infty}\mathrm{d}\chi\int_{0}^{1}\mathrm{d}R\,\chi\left[\mathcal{K}_{i}^{\mathrm{A}}(\chi)\mathcal{K}_{j}^{\mathrm{B}}(R\chi)+\mathcal{K}_{j}^{\mathrm{B}}(\chi)\mathcal{K}_{i}^{\mathrm{A}}(R\chi)\right]w_{\mathrm{AB}}(\chi,R\chi)=\int_{0}^{\infty}\mathrm{d}\chi\int_{0}^{1}\mathrm{d}R\,\chi\tilde{\mathcal{K}}_{ij}^{\mathrm{AB}}(\chi,R\chi)w_{\mathrm{AB}}(\chi,R\chi).$ (44)
@@
$\mathcal{K}_{i}^{\mathrm{A}}(\chi)=\begin{cases}K_{i}^{\mathrm{A}}(\chi)&\text{for clustering}\\ K_{i}^{\mathrm{A}}(\chi)/\chi^{2}&\text{for lensing.}\end{cases}$ (45)
@@ Appendix B Assessing model errors
$\sigma_{ij}^{\mathrm{AB}}(\ell)=\sqrt{\frac{C_{ii}^{\mathrm{AA}}(\ell)C_{jj}^{\mathrm{BB}}(\ell)+(C_{ij}^{\mathrm{AB}}(\ell))^{2}}{f_{\mathrm{sky}}(2\ell+1)}}$ (46)
where we are using $f_{\mathrm{sky}}=0.4$, the sky fraction for LSST. In particular, we can define a metric that is often referred to as Mean Absolute Error Ratio (MAER):
$\mathrm{MAER}^{\mathrm{AB}}_{ij}(\ell)=\frac{|C_{ij}^{\mathrm{AB}}(\ell)-C_{ij,\mathrm{bench}}^{\mathrm{AB}}(\ell)|}{\sigma_{ij}^{\mathrm{AB}}(\ell)}$ (47)